OpenStudy (anonymous):

if p and q are the roots of X^2 + 2px +q-6=0 the value of p equals?

OpenStudy (anonymous):

use the property of sum of roots and product of roots...it is in class 10 syllabus

OpenStudy (anonymous):

yes i used but not getting

OpenStudy (anonymous):

sum=-b/a product= c/a

OpenStudy (anonymous):

ok..what is a,band c here??

OpenStudy (anonymous):

it is ax^2 + bx +c

OpenStudy (anonymous):

so put the values in equation..where are u finding problem??

OpenStudy (anonymous):

p+q=-2p pq=q-6

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

i am stuck at next step

OpenStudy (anonymous):

yes looks so..is q=-19/3

OpenStudy (anonymous):

from pq=q-6 q=q-6/p...

OpenStudy (anonymous):

we have to find the value of p..

OpenStudy (anonymous):

oh it was q=-9 so p=5/3 i think..

OpenStudy (anonymous):

the answer is -1 , 2

OpenStudy (anonymous):

\[x ^{2}+(\alpha + \beta)x + \alpha \beta\]...

OpenStudy (anonymous):

@satellite73 plzz help

OpenStudy (anonymous):

plzz antone solve for it....

jhonyy9 (jhonyy9):

so than p+q=-2p and p*q=q-6 p+q=-2p p*q=q-6 - so just you need to solve this system for q and p

OpenStudy (anonymous):

\[p+q=-2p\] \[pq=q-6\]

OpenStudy (anonymous):

first one gives \(q=-3p\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then too the eq is in p and q

OpenStudy (anonymous):

-3p^2=q-6

OpenStudy (maheshmeghwal9):

who will put q=-3p in r.h.s

OpenStudy (anonymous):

substitute in to second one get \[-3p^2=-3p-6\] \[3p^2-3p-6=0\] \[3(p^2-p-2)=0\] \[(p-2)(p+1)=0\] \[p=2, p=-1\]

OpenStudy (anonymous):

probably easier to solve using quadratic formula or easier still to complete the square

OpenStudy (maheshmeghwal9):

I think u should get it now @shameer1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ignore 9π http://screencast.com/t/MdMY5b0u